# VideoLLM Knows When to Speak: Enhancing Time-Sensitive Video Comprehension with Video-Text Duet Interaction Format

### v1

@@ Abstract
Recent researches on video large language models (VideoLLM) predominantly focus on model architectures and training datasets, leaving the interaction format between the user and the model under-explored. In existing works, users often interact with VideoLLMs by using the entire video and a query as input, after which the model generates a response. This interaction format constrains the application of VideoLLMs in scenarios such as live-streaming comprehension where videos do not end and responses are required in a real-time manner, and also results in unsatisfactory performance on time-sensitive tasks that requires localizing video segments. In this paper, we focus on a video-text duet interaction format. This interaction format is characterized by the continuous playback of the video, and both the user and the model can insert their text messages at any position during the video playback. When a text message ends, the video continues to play, akin to the alternative of two performers in a duet. We construct MMDuetIT, a video-text training dataset designed to adapt VideoLLMs to video-text duet interaction format. We also introduce the Multi-Answer Grounded Video Question Answering (MAGQA) task to benchmark the real-time response ability of VideoLLMs. Trained on MMDuetIT, MMDuet demonstrates that adopting the video-text duet interaction format enables the model to achieve significant improvements in various time-sensitive tasks (76% CIDEr on YouCook2 dense video captioning, 90% mAP on QVHighlights highlight detection and 25% R@0.5 on Charades-STA temporal video grounding) with minimal training efforts, and also enable VideoLLMs to reply in a real-time manner as the video plays. Code, data and demo are available at: https://github.com/yellow-binary-tree/MMDuet.
@@ 1 Introduction
Videos are becoming an increasingly important medium to acquire information on a daily basis. Powered by recent advancements in large language models (LLMs) [28, 11, 25, 2, 37] and vision encoders [22, 39, 26, 21, 30], several video large language models (VideoLLM) [17, 18, 16, 15, 40, 32] have already demonstrated strong abilities for holding conversations and answering questions about videos. A common feature of these models is using visual encoders to encode all frames sampled from the entire video at first, and integrate them into text input by concatenating them to input embeddings or using cross attention.
@@
However, this all-along used whole video interaction has the following three defects, which hinder the performance and real-world usage scenarios of VideoLLMs: Firstly, it does not admit timely interactions. As the video is often input as a whole, this limits its usage in more scenarios like live broadcasts or surveillance videos, in which the video does not end at a specific time. Even if we can segment the video into multiple fixed-length clips for input, the model still cannot generate responses in a real-time manner when necessary. Instead, it must wait until the end of the clip before starting to reply, which negatively impacts the timeliness of the interaction. Secondly, it performs unfavorably on time-sensitive video comprehension tasks. In this paper we use “time-sensitive tasks” to refer to tasks in which the model is required to provide responses that include specific times in the video, such as temporal video grounding [13, 3, 8], video highlight detection [14], dense video captioning [41, 13], grounded video question answering [34], etc. Existing approaches refer to the video timeline by generating text that contains time information, such as second numbers like “3.5 – 7.1 seconds” or timeline percentages like “from 30 to 50”. However, the performance of these models has been consistently unsatisfactory, possibly due to LLMs’ limited ability to handle numbers [24]. Thirdly, it suffers from sub-optimal performance of describing long videos. When generating responses relevant to a specific segment in a lengthy video, the model needs to first implicitly “retrieve” this relevant segment from the entire video that is already encoded in the context, which is challenging for most VideoLLMs [33].
@@
In this work, we formalize the Video-Text Duet Interaction Format, an interaction method that aims to enhance VideoLLMs by addressing the aforementioned issues. An illustration of the whole video interaction format and the video-text duet interaction format is shown in Fig. 1. With our video-text duet interaction format, the video is continuously played and input to the model frame-by-frame. Both the user and model can insert their text messages right after any frame during the video play. When a text message ends, the video continues to play, akin to the show of two performers in a duet. This is implemented by also defining the “video stream” as a participant in the dialogue, with its messages only comprising video frames. When a dialogue turn from either the user or the model ends, the video stream can have the floor and input video frames to the model until another turn is started by either the user or the model. This improves the timeliness of interaction and better suits real-world applications such as live-streaming or surveillance video comprehension. Moreover, by inserting responses to the video where is most relevant, the model can learn to generate responses by referencing a smaller but fine-grained fraction of the video before this position. In this manner, it facilitates information retrieval to describe lengthy videos, as well as enables a response to be “grounded” at the targeted position of the video. We believe this design contributes to addressing the above discussed three issues of existing VideoLLMs.
@@
To prove the effectiveness of the video-text duet interaction format, we construct MMDuetIT, a dataset to facilitate the training of a versatile VideoLLM following the video-text duet interaction format. MMDuetIT is constructed by reformatting existing dense video captioning and temporal video grounding datasets by inserting each turn of the conversation at a proper position within the video. We also propose Multi-Answer Video Grounded QA (MAGQA), a novel task that requires the model to generate answers at appropriate positions within the video in a real-time manner, to align with potential applications of live-streaming video comprehension. We also train MMDuet, a VideoLLM that implements our proposed video-text duet interaction format. Initialized with LLaVA-OneVision [15] and trained with MMDuetIT at a low cost, MMDuet achieves significant performance improvement in various time-sensitive tasks, and is able to generate responses in real-time as the video plays.
@@ 2 Related Works
2.1 Video Large Language Models
The advancement of large language models (LLMs) and visual encoders has led to numerous efforts on their integration, aiming to utilize the powerful understanding and generation abilities of existing LLMs for video-related tasks [17, 18, 16, 15, 32, 36]. These models exhibit a decent ability of video understanding such as captioning or summarizing [36]. However, their performance on time-sensitive tasks is still unsatisfactory, primarily due to two reasons: First, many models are trained on datasets in which video contents are relatively static with annotations not dense enough [40]. It thus constrains the model to discern the temporal differences across various parts of videos. Second, differing from image models that can easily capture word patch alignment [5], most VideoLLMs are not explicitly trained to distinguish differences between different clips in the same video.
2.2 Localizing Video Segments with VideoLLMs
Recent works attempt to empower VideoLLMs with the ability to localize and represent segments in videos, and thus achieve better performance on tasks like temporal video grounding or dense video captioning. These works explore new ways on how to easily represent video clips with texts, such as second numbers of timestamp (TimeChat [23]), timeline percentage (VTimeLLM [10]) or using special textual tokens (VTG-LLM [6], Grounded-VideoLLM [29]). However, their performance has not been satisfactory yet, possibly due to LLMs’ limited ability to accurately count and generate numbers [24] to localize each video frame. To alleviate this issue, HawkEye [31] uses a coarse-grained method by referring to a larger fraction of the video, but it requires multiple rounds of recursive grounding to precisely locate a segment and may not express multiple segments at a time.
The work most similar to our motivation is VideoLLM-Online [1], which proposes a framework named LIVE for training VideoLLMs to interrupt video streams and insert responses. However, they only finetune a model on Ego4D [4] and COIN [27] to demonstrate the LIVE training and inference, and do not explore on how the model capabilities vary with this new type of interaction, especially the zero-shot performance on time-sensitive tasks.
@@ 3 The Video-Text Duet Interaction Format
In Sec. 1, we have defined the concept of “interaction format” with two aspects (i.e., chat template & turn-taking rule), as well as the drawbacks of the commonly-used whole video interaction format. Now we re-emphasize and formalize our video-text duet interaction format, which is completely different from previous to implement VideoLLMs.
(1) For the chat template, inspired by but different from the LIVE framework which is used to implement VideoLLM-Online [1], we consider the video stream as a conversation participant just like the role of user/assistant, and the input sequence consists of alternating turns among these three roles. Take the LLM Qwen2 [37] as an example, the token sequence input to the model should be as follows:
[…]
(2) For the turn-taking rule, when the turn of the user or assistant ends, the video stream can take the floor and start its turn to input video frames. When each single frame is consumed, both the user and the assistant role can interrupt the video stream at any time, and start its own turn to query or generate a response, as totally decided by the user or the assistant, respectively.
@@ 4.1 Model Structure
We propose MMDuet, a model trained following the video-text duet interaction format, which can thus autonomously decide at what position in the video to generate what response. Like almost all existing VideoLLMs, MMDuet consists of three components: 1) a visual encoder that encodes sampled frames from the video to visual feature, 2) a projector that transforms the encoded visual feature to a list of visual tokens that is aligned into the LLM textual embedding space, and 3) a transformer-decoder-based LLM that takes both textual and visual tokens as input and uses its language modeling head to predict the next token. Note that we simply use a linear layer instead of video-level or sliding-window Qformers [17, 23] as the projector, as in the video-text duet interaction format each frame is added to LLM input sequence independently.
The only difference in model structure between our MMDuet and existing VideoLLMs is that we add two more heads in addition to the language modeling head (LM Head) of the LLM, namely the informative head and the relevance head, for determining whether to start a response after each frame. Each head is a linear layer and has a weight with shape $h\times 2$, where $h$ is the hidden size of the used LLM. Each head takes the final layer hidden state of the last visual token of each frame as input, and performs a binary classification. To be specific, 1) the informative head is designed to predict how much new information is acquired upon viewing the current frame. If the model can obtain a “significant amount” of new information upon viewing a new frame (which we will further discuss in Sec. 5.1), it should classify this frame as TRUE category; otherwise, it should classify it as FALSE. 2) The relevance head is designed to predict whether the current frame is related to the user query. Similarly, TRUE category means to be related, while FALSE means not. We denote the probability of TRUE category of informative head and relevance head as informative score and relevance score for each sampled video frame. These two scores will be used to decide whether the model (i.e., assistant role) should interrupt the video and start its own turn. Compared with VideoLLM-Online [1] that makes this decision by predicting a special token using the LM Head, our design of using two additional heads has the following merits: (1) By combining two scores we can flexibly set different criteria for response generation, rather than only relying on the logits of the special tokens; (2) The relevance head can be used to precisely perform temporal video grounding and highlight detection tasks, expanding the application scenarios of MMDuet.
@@ 4.2 Inference Procedure
A python-style pseudocode of the inference process is shown in LABEL:alg:inference. When consuming every single sampled frame of the video, we first check if there is a user query happening at this time. If yes, we first input this user turn to the model. Then the sampled frame is input to the model, after which the informative score and relevance score are calculated. In addition, we use a function need_response to estimate whether the model should generate an assistant response according to the informative scores and relevance scores for this frame along with previous frames. If yes, the generate function of the LLM outputs a response. Different need_response functions can be designed depending on the specific task, which is introduced in the experiment section (Sec. 6). This process can be efficiently implemented by updating the KV Cache each time when a frame or text is input or generated.
# Input:
# system_prompt
# video: list of frames
# fps: frames per second to sample from video
# user_turns: list of (time, text) sorted by time
# Output:
# model_turns: generated list of (time, text)
model_turns = []
v_inf_list, v_rel_list = [], []
kv_cache = model(system_prompt)
time = 0
for frame in video:
if len(user_turns) and time>=user_turns[0].time:
kv_cache = model(kv_cache, user_turns[0].text)
user_turns = user_turns[1:]
kv_cache, v_inf, v_rel = model(kv_cache, frame)
v_inf_list.append(v_inf) # informative score
v_rel_list.append(v_rel) # relevance score
if need_response(v_inf_list, v_rel_list):
kv_cache, response = model.generate(kv_cache)
model_turns.append((time, response))
time += 1 / fps
Listing 1: Inference Process of MMDuet
@@ 5 MMDuetIT: Dataset for Training MMDuet
We build MMDuetIT, a dataset for training the MMDuet model to learn to calculate the informative and relevance scores, and autonomously output replies at any necessary time in the play of the video. MMDuetIT is composed of three different types of tasks that benefit our model training: dense captioning, multi-answer grounded video question answering, and temporal video grounding. An example of the input format for each task is listed in the appendix.

### v2

@@ Abstract
Recent researches on video large language models (VideoLLM) predominantly focus on model architectures and training datasets, leaving the interaction format between the user and the model under-explored. In existing works, users often interact with VideoLLMs by using the entire video and a query as input, after which the model generates a response. This interaction format constrains the application of VideoLLMs in scenarios such as live-streaming comprehension where videos do not end and responses are required in a real-time manner, and also results in unsatisfactory performance on time-sensitive tasks that requires localizing video segments. In this paper, we focus on a video-text duet interaction format. This interaction format is characterized by the continuous playback of the video, and both the user and the model can insert their text messages at any position during the video playback. When a text message ends, the video continues to play, akin to the alternative of two performers in a duet. We construct MMDuetIT, a video-text training dataset designed to adapt VideoLLMs to video-text duet interaction format. We also introduce the Multi-Answer Grounded Video Question Answering (MAGQA) task to benchmark the real-time response ability of VideoLLMs. Trained on MMDuetIT, MMDuet demonstrates that adopting the video-text duet interaction format enables the model to achieve significant improvements in various time-sensitive tasks (76% CIDEr on YouCook2 dense video captioning, 90% mAP on QVHighlights highlight detection and 25% R@0.5 on Charades-STA temporal video grounding) with minimal training efforts, and also enable VideoLLMs to reply in a real-time manner as the video plays.
VideoLLM Knows When to Speak: Enhancing Time-Sensitive Video Comprehension with Video-Text Duet Interaction Format
Yueqian Wang1, Xiaojun Meng2, Yuxuan Wang3, Jianxin Liang1, Jiansheng Wei2, Huishuai Zhang1,4, Dongyan Zhao1,4, 1Wangxuan Institute of Computer Technology, Peking University 2Huawei Noah’s Ark Lab 3Beijing Institute for General Artificial Intelligence 4State Key Laboratory of General Artificial Intelligence Correspondence: zhanghuishuai@pku.edu.cn, zhaodongyan@pku.edu.cn
@@ 1 Introduction
Videos are becoming an increasingly important medium to acquire information on a daily basis. Powered by recent advancements in large language models (LLMs) Touvron2023Llama2O; Jiang2023Mistral7; Shao2024DeepSeekV2AS; Dubey2024Llama3; Yang2024Qwen2TR and vision encoders Radford2021LearningTV; Zhai2023SigmoidLF; Sun2023EVACLIPIT; Oquab2023DINOv2LR; Wang2024InternVideo2SV, several video large language models (VideoLLM) Li2023MVBenchAC; Liu2024STLLMLL; Li2024LLaVANeXTInterleaveTM; Li2024LLaVAOneVisionEV; Zhang2024LLaVAVideo; Wang2024EfficientTE have already demonstrated strong abilities for holding conversations and answering questions about videos. A common feature of these models is using visual encoders to encode all frames sampled from the entire video at first, and integrate them into text input by concatenating them to input embeddings or using cross attention.
@@
However, this all-along used whole video interaction has the following two defects, which hinder the performance and real-world usage scenarios of VideoLLMs: Firstly, it does not admit timely interactions. As the video is often input as a whole, this limits its usage in more scenarios like live broadcasts or surveillance videos, in which the video does not end at a specific time. Even if we can segment the video into multiple fixed-length clips for input, the model still cannot generate responses in a real-time manner when necessary, as it does not know whether it is feasible and appropriate to reply at the end of this clip. Secondly, it performs unfavorably on time-sensitive video comprehension tasks. In this paper we use “time-sensitive tasks” to refer to tasks in which the model is required to provide responses that include specific times in the video, such as temporal video grounding Krishna2017DenseCaptioningEI; Gao2017TALLTA; Hendricks2017LocalizingMI, video highlight detection Lei2021QVHighlightsDM, dense video captioning Zhou2017TowardsAL; Krishna2017DenseCaptioningEI, grounded video question answering Xiao2023NextGQA, etc.
@@
In this work, we formalize the Video-Text Duet Interaction Format, an interaction method that aims to enhance VideoLLMs by addressing the aforementioned issues. An illustration of the whole video interaction format and the video-text duet interaction format is shown in Fig.˜1. With our video-text duet interaction format, the video is continuously played and input to the model frame-by-frame. Both the user and model can insert their text messages right after any frame during the video play. When a dialogue turn from either the user or the model ends, the video stream can have the floor and input video frames to the model until another turn is started by either the user or the model, akin to the show of two performers in a duet. This improves the timeliness of interaction and better suits real-world applications such as live-streaming or surveillance video comprehension. Moreover, by inserting responses to the video where is most relevant, the model can learn to generate responses by referencing a smaller but fine-grained fraction of the video before this position. In this manner, it facilitates information retrieval to describe lengthy videos, as well as enables a response to be “grounded” at the targeted position of the video. We believe this design contributes to addressing the above discussed issues of existing whole video VideoLLMs.
@@
To prove the effectiveness of the video-text duet interaction format, we construct MMDuetIT, a dataset to facilitate the training of a versatile VideoLLM following the video-text duet interaction format. We propose Multi-Answer Video Grounded QA (MAGQA), a novel task that requires the model to generate answers at appropriate timespans in a real-time manner to align with potential applications of live-streaming video comprehension. We also train MMDuet, a VideoLLM that implements our proposed video-text duet interaction format. Initialized with LLaVA-OneVision Li2024LLaVAOneVisionEV and trained with MMDuetIT at a low cost, MMDuet achieves significant performance improvement in various time-sensitive tasks, and is able to generate responses in real-time as the video plays.
@@ 2 Related Works
The advancement of large language models (LLMs) and visual encoders has led to numerous efforts on their integration, aiming to utilize the powerful understanding and generation abilities of existing LLMs for video-related tasks Li2023MVBenchAC; Liu2024STLLMLL; Li2024LLaVANeXTInterleaveTM; Li2024LLaVAOneVisionEV; Wang2024EfficientTE; ijcai2023p582. These models exhibit a decent ability of video understanding such as captioning or summarizing ijcai2023p582. However, their performance on time-sensitive tasks is still unsatisfactory.
Recent works attempt to empower VideoLLMs with the ability to localize and represent segments in videos, and thus achieve better performance on tasks like temporal video grounding or dense video captioning. These works explore new ways on how to easily represent video clips with texts, such as second numbers of timestamp (TimeChat Ren2023TimeChatAT), timeline percentage (VTimeLLM Huang2023VTimeLLMEL) or using special textual tokens (VTG-LLM Guo2024VTGLLMIT, Grounded-VideoLLM Wang2024GroundedVideoLLMSF). However, their performance has not been satisfactory yet, possibly due to LLMs’ limited ability to accurately count and generate numbers Schwartz2024NumeroLogicNE to localize each video frame. To alleviate this issue, HawkEye Wang2024HawkEyeTV uses a coarse-grained method by referring to a larger fraction of the video, but it requires multiple rounds of recursive grounding to precisely locate a segment and may not express multiple segments at a time.
The work most similar to our motivation is VideoLLM-Online Chen2024VideoLLMonlineOV, which proposes a framework named LIVE for training VideoLLMs to interrupt video streams and insert responses. However, they only finetune a model on Ego4D Grauman2021Ego4DAT and COIN Tang2019COIN to demonstrate the LIVE training and inference, and do not explore on how the model capabilities vary with this new type of interaction, especially the zero-shot performance on time-sensitive tasks.
@@ 3 The Video-Text Duet Interaction Format
In Section˜1, we have defined the concept of “interaction format” with two aspects (i.e., chat template & turn-taking rule), as well as the drawbacks of the commonly-used whole video interaction format. Now we re-emphasize and formalize our video-text duet interaction format, which is completely different from previous to implement VideoLLMs.
(1) For the chat template, inspired by but different from the LIVE framework which is used to implement VideoLLM-Online Chen2024VideoLLMonlineOV, we consider the video stream as a conversation participant just like the role of user/assistant, and the input sequence consists of alternating turns among these three roles. (2) For the turn-taking rule, when the turn of the user or assistant ends, the video stream can take the floor and start its turn to input video frames. When each single frame is consumed, both the user and the assistant role can interrupt the video stream at any time, and start its own turn to query or generate a response, as totally decided by the user or the assistant, respectively.
@@ 4.1 Model Structure
We propose MMDuet, a model trained following the video-text duet interaction format, which can thus autonomously decide at what position in the video to generate what response. Like almost all existing VideoLLMs, MMDuet consists of three components: 1) a visual encoder that encodes sampled frames from the video to visual feature, 2) a linear projector that transforms the encoded visual feature to a list of visual tokens that is aligned into the LLM textual embedding space, and 3) a transformer-decoder-based LLM that takes both textual and visual tokens as input and uses its language modeling head to predict the next token.
The only difference in model structure between our MMDuet and existing VideoLLMs is that we add two more heads in addition to the language modeling head (LM Head) of the LLM, namely the informative head and the relevance head, for determining whether to start a response after each frame. Each head is a linear layer and has a weight with shape $h\times 2$, where $h$ is the hidden size of the used LLM. Each head takes the final layer hidden state of the last visual token of each frame as input, and performs a binary classification. To be specific, 1) the informative head is designed to predict how much new information is acquired upon viewing the current frame. If the model can obtain a “significant amount” of new information upon viewing a new frame (which we will further discuss in Section˜5.1), it should classify this frame as TRUE category; otherwise, it should classify it as FALSE. 2) The relevance head is designed to predict whether the current frame is related to the user query. Similarly, TRUE category means to be related, while FALSE means not. We denote the probability of TRUE category of informative head and relevance head as informative score and relevance score for each sampled video frame. These two scores will be used to decide whether the model (i.e., assistant role) should interrupt the video and start its own turn. Compared with VideoLLM-Online Chen2024VideoLLMonlineOV that makes this decision by predicting one special token using the LM Head, our design has the following merits: (1) The ground truth labels of informative scores and related scores are acquired based on the characteristic of the video itself, rather than on ad-hoc response decisions. Therefore, there are better labels for models to converge during training. (2) By combining two scores we can flexibly set different criteria for response generation, rather than only relying on the logits of one special token; (3) The relevance head can be used to precisely perform temporal video grounding and highlight detection tasks, expanding the application scenarios of MMDuet.
@@ 4.2 Inference Procedure
When consuming every single sampled frame of the video, we first check if there is a user query happening at this time. If yes, we first input this user turn to the model. Then the sampled frame is input to the model, after which the informative score and relevance score are calculated. We use a function need_response to estimate whether the model should generate an assistant response according to the informative scores and relevance scores for this frame along with previous frames. If yes, the generate function of the LLM outputs a response. Different need_response functions can be designed depending on the specific task, which is introduced in the experiment section (Section˜6). This process can be efficiently implemented by updating the KV Cache each time when a frame or text is input or generated, and a python-style sudo code is provided in Section˜B.3.
@@ 5 MMDuetIT: Dataset for Training MMDuet
We build MMDuetIT, a dataset for training the MMDuet model to learn to calculate the informative and relevance scores, and autonomously output replies at any necessary time in the play of the video. MMDuetIT is composed of three different types of tasks that benefit our model training: dense captioning, multi-answer grounded video question answering, and temporal video grounding. An example of the input format for each task is listed in Appendix˜D.